Kim is 3 times as old as her cousin. The total of their ages is 36 years. How old is Kim's cousin?
Answer by addingup(3677)

K = 3c
K+c = 36 substitute for K:
3c+c = 36
4c = 36
c = 36/4 Divide to get the cousin's age, then multiply the cousin's age times 3 to get Kim's age.